Plzz give me a clear clarification abt hw to solve it.
Find fog and gof
equation-for-fog-and-gof

Grade:Select Grade

1 Answers

noogler
8 years ago
the common domain of f(x)andg(x) is [-1,2]
 
fog=f(x2)=x2+1         -1
     =f(x2)=2x2+1        1
gof=g(x+1)=(x+1)2        -1
     =g(2x+1)=(2x+1)2      1